# A Comprehensive Survey on Retrieval Methods in Recommender Systems

### v1

@@ 4.6. Multi-Interest Methods
The deep neural retrieval methods mentioned above typically assign a single vector to represent each user, encapsulating their preferences and behaviors. However, this single vector representation often falls short in capturing the complexity and diversity of a user’s multiple interests. As a result, recent works have developed models that encapsulate multiple interest vectors for each user. One of the pioneering models is MIND (Li et al., 2019c), which introduces a dynamic routing mechanism to assign each user interaction to specific interests. ComiRec (Cen et al., 2020) builds on this by employing a self-attentive mechanism to distill user interests from their activity patterns. Further advancements include models like SINE (Tan et al., 2021) and Octopus (Liu et al., 2020a), which establish interest pools and use attention mechanisms to activate relevant interests based on historical interactions. DGCF (Wang et al., 2020a) extends these concepts by integrating a dynamic routing mechanism into graph structures, focusing on the independence among interests for enhanced multi-interest learning.
@@ 5.1. Whole Vault Calculation
In the simplest form of retrieval, there is no indexing mechanism, and the system performs whole vault calculations. This approach involves computing the similarity between the user and every item in the candidate pool. While this method can be straightforward to implement, it is computationally expensive and impractical for large-scale systems with millions of items due to the sheer volume of calculations required.
@@ 5.2. Approximate Nearest Neighbor Search
Various techniques and tools facilitate ANN search, such as Faiss (Johnson et al., 2019), Milvus (Wang et al., 2021), and ScaNN (Guo et al., 2020a), which are instrumental in industry applications for efficiently retrieving the top-$K$ nearest vectors. The construction of popular ANN indices, such as HNSW (Malkov and Yashunin, 2018) and IVFPQ in Faiss (Johnson et al., 2019), involves time-consuming procedures, including clustering algorithms like $K$-means. Despite these challenges, the advantages of using ANN search in practical applications, particularly in speeding up the retrieval process and reducing computational loads, make it a valuable technique in modern retrieval systems.
@@ 5.3. Tree Based Indexing
Tree based indexing methods have gained attention in recent years for their ability to efficiently organize and retrieve large-scale items. These methods leverage hierarchical structures to capture the diversity of user interests and represent information more effectively. Tree based indexing methods offer several advantages over vector based approaches. They can significantly reduce the computational complexity by structuring the data hierarchically, enabling logarithmic scale computation relative to the size of the corpus during prediction. Additionally, they allow for the interaction between candidate items and users within the model, enhancing the expressive power of the retrieval method. These methods not only make the use of arbitrary advanced models possible, but also excel at discovering novel and effective recommendations by exploring the entire dataset and utilizing sophisticated deep models to uncover potential user interests. This capability represents a significant leap forward, offering a powerful solution for improving the performance of retrieval.
TDM (Zhu et al., 2018) is a pioneering work in this area. It ties each item in the corpus to a leaf node via clustering, using a tree structure as an index. The model computes the probability value that corresponds to each embedding vector represented by a node in the tree. The user’s interest in a specific node is described by its probability value, while the diversity of user interest is represented by the entire tree. This transforms the retrieval process into a sequence of hierarchical classification tasks, achieving efficient and scalable retrieval. JTM (Zhu et al., 2019) improves upon TDM by synchronizing the objectives of index learning and model training, addressing inconsistencies, and enhancing the model’s overall performance. OTM (Zhuo et al., 2020) further refines this approach by addressing the discrepancy between training and testing in tree based methods. It introduces the notion of bayes optimality under beam search and calibration under beam search, providing general tools for analyzing and improving tree based methods.
@@ 6.2.3. Pairwise Loss
6.3. Metrics
In this section, we introduce some commonly-used evaluation metrics in retrieval methods. Recall@K, Precision@K, and F1@K are closely related metrics in the retrieval stage. We categorize the ground truth labels as either positive or negative, within the predicted labels, this results in classifications such as True Positive (TP), True Negative (TN), False Positive (FP), and False Negative (FN). Recall and Precision indicate the proportion of correctly predicted positive samples relative to the predicted positives and ground truth positives, respectively. The F1-score combines recall and precision into a single metric, as is shown in  equation 13.
(13) $Recall=\frac{TP}{TP+FN}\quad Precision=\frac{TP}{TP+FP}\quad F1=\frac{2*\text{precision }*\text{ recall }}{\text{ precision }+\text{ recall }}$
@@ 7.1. Datasets
To provide a clear understanding of the performance of various retrieval methods, we present the results of several retrieval methods on the MovieLens-1M, Amazon-Book, and Yelp2018 datasets below. Following previous research (He and McAuley, 2016; He et al., 2017), we use the 10-core setting to ensure that each user and item have at least ten interactions. The statistics of these datasets are shown in Table 1.
Table 1. Statistics of the datasets.
Dataset # Users # Items # Samples Density
MovieLens-1M $6,040$ $3,706$ $1,000,209$ 0.04468
Amazon-Book $52,643$ $91,599$ $2,894,108$ 0.00062
Yelp2018 $31,668$ $38,048$ $1,561,406$ 0.00130
@@ 7.2. Experiment Settings
Evaluation Metrics. We adopt two commonly used metrics to evaluate the performance of various retrieval methods: HitRate@K and Recall@K. Specifically, we set K=50 on MovieLens-1M, and K=100 on Amazon-Book and Yelp2018 datasets. See Section 6.3 to have a deeper understanding of different metrics.
@@
Implementation Details. As with previous work (He et al., 2017), for each dataset, we randomly select 80% of each user’s historical interactions for the training set and use the remaining 20% as the test set. Additionally, we randomly select 10% of the interactions from the training set to serve as the validation set. We implement the baseline methods using the RecBole open-source framework and its provided code (Zhao et al., 2021).
@@ 8. CASE STUDY OF CURRENT INDUSTRIAL PRACTICES
In this section, we present a detailed case study of the retrieval pipeline at China Merchants Bank (CMB), a mainstream bank company. We focus on CMB Life APP for credit card [^1] https://english.cmbchina.com/CreditCard/, which attracts millions of daily active users and generates billions of user logs each day through implicit feedback, such as click behaviors. By concentrating on high-frequency consumption scenarios, CMB Life APP has enriched its online service ecosystem. It integrates various scenarios, such as mall products, meal and movie tickets, live broadcasts, and advertisements. Each scenario serves different business units with distinct performance metrics, such as click-through rates and conversion rates.
@@ 8.2. Multi-Channel Retrieval
For Word2Vec, we feed comprehensive text fields from within the scenario such as item titles, attributes, and categories to generate item embeddings, which are then stored in an embedding database. The underlying principle is that items with similar titles, attributes, and categories are more likely to appeal to users. For Node2Vec, we create positive sample pairs from user behavior sequences by using a sliding window of length 2 to generate positive item pair item $i$ and item $j$. Simultaneously, we form negative item pairs by pairing item $i$ with a random item $k$ from the item pool, ensuring that $k\neq j$, to train the Node2Vec retrieval model. In the setting of I2I, we directly store the generated item embeddings, while in the setting of U2I, we obtain user embeddings by weighting the embeddings of the recent sequence of items interacted with by the user. We then use the user embedding to retrieve the top-$K$ items, storing them in the format of $\xrightarrow{}$temList. The rationale is that items interacted with consecutively in a user’s behavior sequence are perceived as more similar by the user, making it likely that they will engage with other items close to item $i$ after visiting it. For DSSM, we generate positive user-item pairs from user behavior sequences and create negative samples by pairing the user with a randomly sampled item from the item pool. These data are then used to train the DSSM retrieval model, which yields item embeddings that are then stored in the embedding database. DSSM may offer more personalization than Node2Vec in this scenario, as the implicit graph structure includes both users and items.
@@
Figure 6 displays the distribution and online performance of the seven retrieval channels mentioned above. In the retrieval stage, the focus is on covering potential user interests with various approaches, making the integration of multiple retrieval results a significant topic. Currently, in the scenario of CMB Life APP for credit cards, our retrieval strategy is divided into two layers. The first layer includes multiple retrieval methods, consisting of three I2I retrieval methods, one U2I retrieval method, and a retrieval method for new and niche items. The second layer involves supplementary retrieval approaches, including popular item retrieval and user tag based popular item retrieval. This supplementary layer is only activated if the first layer’s results are insufficient. For example, if we aim to retrieve 300 items and the first layer retrieves only 180, the second layer will add the top 120 to reach 300. If the first layer already retrieves 300 items, the second layer will not be triggered. As seen in Figure 6, supplementary retrieval accounts for about 50% in this scenario, largely due to the presence of many inactive users.
@@ 9.2. From Single-Stage Retrieval Optimization to Cascade System Studies
In recent years, there has been a notable shift from single-stage retrieval models to exploring multi-stage cascade ranking systems. For instance, Gallagher et al. (Gallagher et al., 2019) explored deriving gradients for cascade rankers, employing end-to-end methods for optimization, while Fan et al. (Fan et al., 2019) advocated for unifying multiple stages into a single stage using hard negative sampling. Furthermore, Fei et al. (Fei et al., 2021) suggested a strategy to share features across different stages. Hron et al. (Hron et al., 2021) suggests using multiple channels together during the retrieval stage and learning how to merge the items retrieved from various channels. Rankflow (Qin et al., 2022) suggest training rankers on their specific data distributions to alleviate the SSB issue and better utilize the interactions between each stage. In CoRR (Huang et al., 2023), the retriever is enhanced by deriving high-quality training signals from the ranker, while the ranker benefits from learning to discriminate hard negatives sampled by the retriever. GPRP (Zheng et al., 2024) addresses both the selection bias in each stage of the system pipeline and the underlying interests of users. The core idea is to estimate the selection bias in subsequent stages and then develop a ranking model that aligns with this bias, ensuring the top-ranked results are delivered in the final output. RAR (Huang et al., 2024) stands for Recall Augmented Ranking, which optimizes the recommendation process by utilizing multi-stage information. ARF (Wang et al., 2023d) introduces a novel perspective on optimizing cascade ranking systems by emphasizing the adaptability of optimization targets to data complexities and model capabilities. These studies highlight the importance of viewing the retrieval process from a cascade system perspective, where each stage is optimized not in isolation but in concert with the others to improve the overall performance and effectiveness of the recommender system.

### v2

@@ 4.6. Multi-Interest Methods
The deep neural retrieval methods mentioned above typically assign a single vector to represent each user, encapsulating their preferences and behaviors. However, this single vector representation often falls short in capturing the complexity and diversity of a user’s multiple interests. As a result, recent works have developed models that encapsulate multiple interest vectors for each user. Let $\mathcal{H}(u)$ be the set of items with which user $u$ has interacted, and let each item $i$ have an embedding $\mathbf{e}_{i}$. A general approach is to generate $K$ interest vectors for user $u$:
(16) $\bigl\{\mathbf{h}_{u,1},\mathbf{h}_{u,2},\dots,\mathbf{h}_{u,K}\bigr\},\qquad\mathbf{h}_{u,k}\;=\;f_{k}\bigl(\{\mathbf{e}_{i}:i\in\mathcal{H}(u)\}\bigr),$
where each $f_{k}(\cdot)$ denotes a mechanism, such as dynamic routing or self-attention, that extracts the $k^{th}$ interest representation from the user’s historical interactions. Consequently, each $\mathbf{h}_{u,k}$ corresponds to one latent interest of user $u$. One of the pioneering models is MIND (Li et al., 2019c), which introduces a dynamic routing mechanism to assign each user interaction to specific interests. ComiRec (Cen et al., 2020) builds on this by employing a self-attentive mechanism to distill user interests from their activity patterns. Further advancements include models like SINE (Tan et al., 2021) and Octopus (Liu et al., 2020a), which establish interest pools and use attention mechanisms to activate relevant interests based on historical interactions. DGCF (Wang et al., 2020a) extends these concepts by integrating a dynamic routing mechanism into graph structures, focusing on the independence among interests for enhanced multi-interest learning.
@@ 5.1. Whole Vault Calculation
In the simplest form of retrieval, there is no indexing mechanism, and the system performs whole vault calculations. This approach involves computing the similarity between the user and every item in the candidate pool.
(17) $s(u,i)=\left\langle\mathbf{h}_{u},\mathbf{h}_{i}\right\rangle\quad\forall i\in\mathcal{I}.$
In  equation 17, $\mathbf{h}_{u}$ and $\mathbf{h}_{i}$ are the embedding vectors for user $u$ and item $i$, and $\langle\mathbf{h}_{u},\mathbf{h}_{i}\rangle$ denotes their inner product. While this method can be straightforward to implement, it is computationally expensive and impractical for large-scale systems with millions of items due to the sheer volume of calculations required.
@@ 5.2. Approximate Nearest Neighbor Search
(18) $s(u,i)=g\left(\mathbf{h}_{u},\mathcal{Q}\left(\mathbf{h}_{i}\right)\right).$
$\mathcal{Q}(\mathbf{h}_{i})$ represents the index-based, compressed or quantized version of $\mathbf{h}_{i}$, and $g(\cdot,\cdot)$ is a scoring function that approximates the similarity between $\mathbf{h}_{u}$ and the indexed representation of $\mathbf{h}_{i}$. Various techniques and tools facilitate ANN search, such as Faiss (Johnson et al., 2019), Milvus (Wang et al., 2021), and ScaNN (Guo et al., 2020a), which are instrumental in industry applications for efficiently retrieving the top-$K$ nearest vectors. The construction of popular ANN indices, such as HNSW (Malkov and Yashunin, 2018) and IVFPQ in Faiss (Johnson et al., 2019), involves time-consuming procedures, including clustering algorithms like $K$-means. Despite these challenges, the advantages of using ANN search in practical applications, particularly in speeding up the retrieval process and reducing computational loads, make it a valuable technique in modern retrieval systems.
@@ 5.3. Tree Based Indexing
Tree based indexing methods have gained attention in recent years for their ability to efficiently organize and retrieve large-scale items. These methods leverage hierarchical structures to capture the diversity of user interests and represent information more effectively. Tree based indexing methods offer several advantages over vector based approaches. They can significantly reduce the computational complexity by structuring the data hierarchically, enabling logarithmic scale computation relative to the size of the corpus during prediction. Additionally, they allow for the interaction between candidate items and users within the model, enhancing the expressive power of the retrieval method. These methods not only make the use of arbitrary advanced models possible, but also excel at discovering novel and effective recommendations by exploring the entire dataset and utilizing sophisticated deep models to uncover potential user interests.
TDM (Zhu et al., 2018) is a pioneering work in this area. It ties each item in the corpus to a leaf node via clustering, using a tree structure as an index. To model user preference, TDM assigns to each non‐leaf node $n$ at level $l$ a score based on its most promising child at level $l+1$. Formally, if we denote by $p^{(l)}(n\mid u)$ the preference of user $u$ for node $n$ at level $l$, then
(19) $p^{(l)}(n\mid u)\;=\;\frac{1}{\alpha^{(l)}}\;\max_{\,c\,\in\,\mathrm{Children}(n)}\;p^{(l+1)}(c\mid u)\,$,
where $p^{(l+1)}(c\mid u)$ is the preference for child node $c$ in the next finer level, and $\alpha^{(l)}$ is a level‐specific normalization constant. In this way, a node’s score propagates upward from the highest‐scoring child, ensuring that the top-$K$ candidates at level $l$ are drawn from among the children of the top-$K$ nodes at level l1l-1. The user’s overall interest is thus represented by the distribution of scores across the entire tree. At inference time, retrieving the top-$K$ leaf items is equivalent to performing a sequence of small classification tasks—selecting the best nodes level by level—yielding an efficient, scalable retrieval process. JTM (Zhu et al., 2019) improves upon TDM by synchronizing the objectives of index learning and model training, addressing inconsistencies, and enhancing the model’s overall performance. OTM (Zhuo et al., 2020) further refines this approach by addressing the discrepancy between training and testing in tree based methods. It introduces the notion of bayes optimality under beam search and calibration under beam search, providing general tools for analyzing and improving tree based methods.
@@ 6.2.3. Pairwise Loss
7. EVALUATION METRICS
In this section, we collect and summarize the most commonly used metrics for assessing retrieval-stage performance. We split these into (1) offline metrics, which are computed on held-out logs or public datasets, and (2) online (A/B) metrics, which are measured in a live serving environment.
7.1. Offline Metrics
Offline evaluation is typically performed on historical logs or public benchmarks. These metrics measure how well a retrieval method ranks known positive interactions versus held-out negatives or unobserved items. Recall@K, Precision@K, and F1@K are closely related metrics in the retrieval stage. We categorize the ground truth labels as either positive or negative, within the predicted labels, this results in classifications such as True Positive (TP), True Negative (TN), False Positive (FP), and False Negative (FN). Recall and Precision indicate the proportion of correctly predicted positive samples relative to the predicted positives and ground truth positives, respectively. The F1-score combines recall and precision into a single metric, as is shown in  equation 25.
[…]
@@ 7.1. Offline Metrics
7.2. Online Metrics
Online evaluation is conducted via A/B testing on live traffic. The primary metrics include click-through rate (CTR), conversion rate (CVR), revenue per mille (RPM), and gross merchandise value (GMV), which are defined in  equation 28, where #\# denotes the quantity:
(28) $CTR=\frac{\#clicks}{\#impressions}\quad CVR=\frac{\#conversions}{\#clicks}\quad RPM=\frac{Revenue}{\#impressions}\times 1000\quad GMV=\sum_{i\in\mathcal{S}}price_{i}\times quantity_{i}.$
When measuring CTR, an ‘impression’ is recorded when an item is rendered in a user’s viewport for at least one second, and a ‘click’ is recorded when the user taps or clicks that item. Closely related to CTR is CVR, which measures the proportion of clicks that result in a downstream action (e.g., ‘Add-to-Cart’ or ‘Purchase’). Since CVR reflects actual revenue-generating behavior, it typically serves as the primary business key performance indicator once engagement (CTR) reaches satisfactory levels. RPM indicates revenue generated per thousand impressions. GMV represents the total monetary value of all transactions originating from recommended items, where $\mathcal{S}$ denotes the set of items purchased.
@@ 8. DATASETS AND EXPERIMENTS
Table 1. Statistics of the datasets.
Dataset # Users # Items # Samples Density
MovieLens-1M $6,040$ $3,706$ $1,000,209$ 0.04468
Amazon-Book $52,643$ $91,599$ $2,894,108$ 0.00062
Yelp2018 $31,668$ $38,048$ $1,561,406$ 0.00130
@@ 8.1. Datasets
To provide a clear understanding of the performance of various retrieval methods, we present the results of several retrieval methods on the MovieLens-1M, Amazon-Book, and Yelp2018 datasets below. For Amazon-Book and Yelp2018, following previous work (He and McAuley, 2016; He et al., 2017), we apply the 10-core filtering to ensure data quality, retaining only users and items with at least ten interactions. Due to its relatively smaller size, MovieLens-1M is used in its original form without core filtering. The statistics of these datasets are shown in Table 1.
@@ 8.2. Experiment Settings
Evaluation Metrics. We adopt two commonly used metrics to evaluate the performance of various retrieval methods: HitRate@K and Recall@K. Specifically, we set K=50 on MovieLens-1M, and K=100 on Amazon-Book and Yelp2018 datasets. See Section 7 to have a deeper understanding of different metrics.
@@
Implementation Details. As with previous work (He et al., 2017), for each dataset, we randomly select 80% of each user’s historical interactions for the training set and use the remaining 20% as the test set. Additionally, we randomly select 10% of the interactions from the training set to serve as the validation set. We implement these methods using the RecBole open-source framework [^1] https://github.com/RUCAIBox/RecBole and its provided code (Zhao et al., 2021), and we adopt the optimal hyperparameters listed on the RecBole website. [^2] https://recbole.io/hyperparameters/index.html
@@ 9. CASE STUDY OF CURRENT INDUSTRIAL PRACTICES
In this section, we present a detailed case study of the retrieval pipeline at China Merchants Bank (CMB), a mainstream bank company. We focus on CMB Life APP for credit card [^3] https://english.cmbchina.com/CreditCard/, which attracts millions of daily active users and generates billions of user logs each day through implicit feedback, such as click behaviors. CMB Life APP has enriched its online service ecosystem by integrating various scenarios. In the meal-and-movie ticketing service, users receive recommendations for merchants offering personalized dining and cinema discounts. In shopping malls, consumers are presented with products redeemable through accumulated loyalty points. On homepage feeds, trending news items appear based on users’ interests. In automotive marketplaces, platforms recommend car brands alongside customized down-payment and monthly-payment plans tailored to individual financial profiles. Each scenario serves different business units with distinct performance metrics, such as click-through rates and conversion rates.
@@ 9.2. Multi-Channel Retrieval
[…]
@@
Figure 7 displays the distribution and online performance of the seven retrieval channels mentioned above. It is important to note that this section evaluates the live production system, diverging from the offline benchmark analysis in Section 8.3; as the experimental contexts and objectives differ, the results from these two sections are not directly comparable. In the retrieval stage, the focus is on covering potential user interests with various approaches, making the integration of multiple retrieval results a significant topic. Currently, in the scenario of CMB Life APP for credit cards, our retrieval strategy is divided into two layers. The first layer includes multiple retrieval methods, consisting of three I2I retrieval methods, one U2I retrieval method, and a retrieval method for new and niche items. The second layer involves supplementary retrieval approaches, including popular item retrieval and user tag based popular item retrieval. This supplementary layer is only activated if the first layer’s results are insufficient. For example, if we aim to retrieve 300 items and the first layer retrieves only 180, the second layer will add the top 120 to reach 300. If the first layer already retrieves 300 items, the second layer will not be triggered. As seen in Figure 7, supplementary retrieval accounts for about 50% in this scenario, largely due to the presence of many inactive users.
@@ 10.2. From Single-Stage Retrieval Optimization to Cascade System Studies
In recent years, there has been a notable shift from single-stage retrieval models to exploring multi-stage cascade ranking systems (Huang et al., 2023; Wang et al., 2023d; Huang et al., 2024; Zhu et al., 2023). For instance, Gallagher et al. (Gallagher et al., 2019) explore deriving gradients for cascade rankers, employing end-to-end methods for optimization, while Fan et al. (Fan et al., 2019) advocate for unifying multiple stages into a single stage using hard negative sampling. Furthermore, Fei et al. (Fei et al., 2021) suggest a strategy to share features across different stages. Hron et al. (Hron et al., 2021) suggest using multiple channels together during the retrieval stage and learning how to merge the items retrieved from various channels. Rankflow (Qin et al., 2022) suggest training rankers on their specific data distributions to alleviate the SSB issue and better utilize the interactions between each stage. In CoRR (Huang et al., 2023), the retriever is enhanced by deriving high-quality training signals from the ranker, while the ranker benefits from learning to discriminate hard negatives sampled by the retriever. GPRP (Zheng et al., 2024) addresses both the selection bias in each stage of the system pipeline and the underlying interests of users. The core idea is to estimate the selection bias in subsequent stages and then develop a ranking model that aligns with this bias, ensuring the top-ranked results are delivered in the final output. RAR (Huang et al., 2024) stands for Recall Augmented Ranking, which optimizes the recommendation process by utilizing multi-stage information. ARF (Wang et al., 2023d) introduces a novel perspective on optimizing cascade ranking systems by emphasizing the adaptability of optimization targets to data complexities and model capabilities. These studies highlight the importance of viewing the retrieval process from a cascade system perspective, where each stage is optimized not in isolation but in concert with the others to improve the overall performance and effectiveness of the recommender system.
@@ 11. CONCLUSION
ACKNOWLEDGMENTS
This work is supported by China Merchants Bank Credit Card Center. The Shanghai Jiao Tong University team is partially supported by National Natural Science Foundation of China (62322603, 62177033, 624B2096).